Privacy and Security Notice

FEL Upgrade Project Weekly Brief
September 17-21, 2001

 
Highlights:
We devoted a full week of FEL operations to the second run for laser material damage studies funded by the
Joint Technology Office (JTO).

Management:
Four proposals were completed and submitted to the DoD JTO on Monday, Sept. 17th.  These proposals are
concerned with technology development that can help push the power of the FEL upgrade or improve the
cost/reliability of the control and power systems.  Proposals were submitted on the topics:  (1) development of
cryogenic mirrors for high power loading; (2) development of drive laser components for increased
performance/reliability; and development of control/power components for (3) diagnostics and (4) magnet trim power supplies based on the commercially available CAN bus technology (used in the auto industry).

We reviewed the technical proposal that we received from the Aerospace Corporation concerning our request
for the design and construction of UV laser processing tools to be used with the UV FEL.  Aerospace Corp. was very responsive to our request and we look forward to working with our colleagues at Aerospace Corp. in the execution of the planned subcontract.

We have received a request from the IDA for a briefing and site visit on the Jefferson Lab FEL program for later in the month.

We completed final preparations for next week's DOE/SURA annual review of Jefferson Lab's Science and
Technology.  Two reviewers from the accelerator physics community will be paying attention to the FEL
program.

A proposal to the state of Virginia's Commonwealth Technology Research Fund was completed and submitted
through the primary sponsoring institution, the College of William and Mary for support of the installation and
re-commissioning of the Helios x-ray source in a planned addition to the FEL Facility.

WBS 4 (Injector):
Finished assembly of the cathode shield test assembly and put it under vacuum. The ball electrode rough cut had the interior electro-polished. Port tubes for the gun chamber are being electro-polished. Design/Detail work
continued on the rear portion of the gun.

WBS 5 (SRF):
Space frame and vacuum vessel for the FEL cryomodule is scheduled to be completed in 2 weeks.

WBS 6 (RF):
Zone 3 - RF testing continues with 6 kW having been generated by position 1.  The other positions will be
powered today and early next week.  Several cables in the vault will be checked over the weekend when the
vault is open.

Zone 4 - The missing equipment has started to return.  Three RF Control Modules have "found" their way back
home.  The balance will be returned in a few weeks.

Quarter - All major HVPS power  components have been ordered and many have been received.  The 50kW
RF Loads will be returned to the vendor in December to upgrade them to 100 kW.  The First Article Tests were
successfully conducted on the first 100 kW klystron.  The test results are being reviewed.

Gun HVPS - The specifications for the HVPS Tank have been firmed and the actual design is starting.

WBS 8 (Instrumentation):
Designs have been completed for a Digital Isolation Module and a TTL version Isolation Module.  The
Video/Line Driver module was debugged and is working.  The Rabbit board is being used to control a filter
wheel.  This is the second application of this particular controller.  The clean area for beam viewer assembly is
complete.  Video was hard-wired into the User Lab 3 patch panel completing the installation of that panel.
Testing of the Optical Beam Position Monitors continues.

Additional time was spent supporting operations for the JTO run and concurrent nano-tube run.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  Discussions were held on tolerance issues and an amendment sent to the offerors that will ask for best and
    final offers.
Injector Dipoles (DU/DV)
o  The Large Magnetís (GU) details are complete and the final round of rework on the Assembly Drawing is
    nearly complete.
Arc Dipoles (GY, GX, GQ)
o  AES has completed their manipulation of the magnetic model of the Reverse Bend  (GQ).  They achieved
    uniform magnetic field over 80 % of the axial volume add a taper to get uniform field and integral.  They
    adjusted the gap height and faces to get the correct bulk field and effective length.
o  Design continued on backchecking the detail drawings of all three magnets with the comments  and
    corrections.
o  They are almost complete with the rework of details of the 180° Dipole (DY) and the assembly drawings,
    incorporating the imperatives of the magnetic model and the suggestions that came from checking.
Quadrupoles
3 inch quad (QX)
o  The assembly and the assembly drawing are in the procurement cycle.
o  Work continues on cores and coils with a batch of 64 coils ready to ship.
Trim Quad (QT)
o This magnet is in the procurement cycle.
Sextupole (SF)
o  Work continued on the magnetic model as well as the drawings.
Octupole (OT)
o  DULY Research had excellent results from their magnetic modeling.  They achieved less
    than 1 % error field (The specification) using a simple truncated mushroom cap pole of a
    mathematically defined shape and ring-like, non-connected field clamp.  They were able to
    use the coils in the model that we will recycle from the surplus Injector Correctors from the
    CEBAF machine.
o  They are working on a plan for the drafting of the magnet.
Corrector Dipoles (DB, DJ)
o  The first, very good looking coil was received from Milhous Co. and they were given the go
    ahead to manufacture all the coils.
Beam Line and Vacuum
o  Work is winding down on the definition of the Optical Chicane chambers.
o  The bid package drawing set for the girders in the return leg was signed off.
o  The recirculation dump concept was finalized by Dave Douglas to the satisfaction of George Neil.  It will
    utilize two Trim Quads (QTs).
o  Work continued on the injector line stands and vacuum chambers.  The chambers were narrowed to fit
    through the vertical portion of the field clamps that DULY found was necessary to obtain the field integral
    uniformity.

WBS 11 (Optics):
We held an internal review of the optical cavity late last Friday (9/14).  Some good comments came out of this
review, and will be incorporated into the design.  Everyone agreed it was a good design, and that we should
proceed with component procurements.  This has been occurring, for example, with orders signed for vacuum
rotary feedthroughs.

Last week we received the first article of our new, near-Brewster windows that also serve as the vacuum/air
interface in the user labs.  These windows incorporate a larger window that permits us to image the entire beam
for profiling, position monitoring, or power.  It is being checked for correct angle, and will then be cleaned for
vacuum use and installed on the optical transport.

Operationally, we have been running experiments for the JTO-funded materials studies.  It has been very
successful, with our enhanced diagnostics.  As of this writing (11:40 am 9/21), we have run 144 targets.  We
have also supported two visiting scientists from NRL, who are using their diagnostics to observe the targets
during irradiation.

Operations/Commissioning:
The schedule for the FEL Fall 2001 run was completed and is being sent to all of the Users and placed on the
Web.